# in Language Models

### v1

@@ 1 Introduction
Large language models (LLMs) continue to grow in size, driven by “scaling laws” that suggest larger models tend to yield better performance (Hestness et al., 2017; Hoffmann et al., 2022; Henighan et al., 2020). Adding more attention blocks increases model capacity, but self-attention is also the most expensive operation in LLMs. Unlike MLP layers, attention incurs a quadratic computational cost with respect to sequence length, making it a dominant factor in inference-time efficiency. However, not all attention blocks are equally important for every task, and some may carry redundant information. In this paper, we explore whether entire attention blocks can be skipped without significant performance degradation.
@@
We evaluate Hopscotch on instruction-tuned models, including Llama-3.1-8B-Instruct and Qwen2.5-7B-Instruct, and find that it can successfully skip up to 7 attention blocks with less than a 3% average performance drop across diverse tasks while yielding up to 15% inference speedup (see Section 4.6). With 4 blocks removed, we retain over 98.6% of baseline accuracy. These results highlight structural redundancy in attention layers. We measure distributional shift in hidden representations and find that Hopscotch significantly reduces deviation from the original model, compared to unscaled attention block skipping. Finally, we show that Hopscotch is compatible with leading quantization techniques, such as GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2024).
@@ 2 Related Work
Skipping computation in LLMs. Several recent studies have explored the possibility of reducing computation in LLMs by selectively skipping parts of the model. Shukor and Cord (2024) show that in multimodal LLMs, certain attention blocks contribute less to performance in vision-language tasks and can be skipped in these cases with minimal degradation. Further investigations on the true impact of attention blocks at different depths (Ben-Artzy and Schwartz, 2024) find that the effect is highly block-specific. Our work builds on these insights by providing a principled mechanism to identify and compensate for less informative attention blocks using learned scaling.
@@ 3 The Hopscotch method
Benchmark Baseline 7 Blocks 4 Blocks
Value Recovery Value Recovery
gsm8k FE 83.10 79.38 95.52% 81.05 97.53%
gsm8k SM 66.94 75.82 113.27% 77.03 115.07%
arc_challenge 64.33 61.17 95.09% 62.42 97.03%
gpqa 29.07 29.39 101.10% 29.53 101.58%
social_iqa 49.85 48.93 98.15% 49.49 99.28%
pawsx 63.26 60.44 95.54% 61.90 97.85%
Table 1: Llama-3.1-8B-Instruct performance on benchmarks after skipping 7 and 4 attention blocks via Hopscotch.
@@ 4.1 Benchmarks and Setup
Figure 2: A plot of the single-epoch average loss per layer removed during step one of Hopscotch for Llama.
Figure 1 compares the benchmark performance of our full method, Hopscotch layer selection with scaling, against four baselines: (i) the original unmodified model (baseline), (ii) random layer removal with no scaling, (iii) random layer removal with Hopscotch scaling, and (iv) Hopscotch-selected layers with no scaling. When simply removing sets of arbitrary attention blocks near the center of the model, performance remained within a few points for all but GSM8K, which degraded drastically. When using Hopscotch scaling, even without layer selection, we were able to recover a significant portion of the GSM8K performance, while also seeing slight improvement in other domains. For this reason, we use the math domain and GSM8K as our primary test case and training recovery set for the Hopscotch method, while continuing to evaluate on the full set of benchmarks.
@@
Table 2: Qwen2.5-7B-Instruct performance after skipping 4 attention blocks via Hopscotch.
@@ 4.5 Compatibility with Quantization
Baseline (Instruct) 0.670 0.839
GPTQ 0.680 0.839
GPTQ + Hopscotch (4) 0.773 0.794
GPTQ + Hopscotch (7) 0.751 0.785
Hopscotch (4) + GPTQ 0.779 0.792
Hopscotch (7) + GPTQ 0.738 0.785
AWQ 0.650 0.812
AWQ + Hopscotch (4) 0.714 0.751
AWQ + Hopscotch (7) 0.695 0.738
Hopscotch (4) + AWQ 0.752 0.783
Hopscotch (7) + AWQ 0.731 0.792
@@ 4.6 Effects on Model Efficiency
In a given forward pass for Llama-3.1-8B-Instruct, we experimentally find that $\sim$66of time is spent in decoder attention blocks. In a model with 32 hidden layers, this means $\sim$2.06of the forward pass is spent in each block. For example, for a query with an average forward pass time of 0.054 seconds, the time spent in a single attention block is on average 0.0011 seconds. The effect of removing a subset of attention blocks on overall inference time is summarized in Table 5.
Attn. Blocks Removed Inference Time Reduction
@@
Given that Llama-3.1-8B-Instruct has 32 decoder layers and a total of $\sim$8B parameters, each layer contains approximately $250$M parameters. Removing one attention block saves roughly $83$M parameters, which is about $1.04\%$ of the model. Removing seven attention blocks yields a total reduction of $7\times 83$M = $581$M parameters, or about $7.28\%$ of the total model size. Since each parameter occupies 2 bytes in float16, we can estimate the memory savings accordingly (Table 6).
Attn. Blocks Removed Parameters Reduction Memory Reduction
@@
Further memory savings will be observed in practice due to reductions in optimizer states during training (e.g., Adam requires 8–12 bytes per parameter, leading to approximately 3× the model size), activation memory for forward and backward passes, gradient storage which is typically the same size as the model, and the KV cache used during inference with long contexts, which can consume 2–4 GB per 1,000 tokens for large models. This also means batch sizes can be potentially increased, further boosting inference gains.
@@ A.1 Data and Loss Function Selection
Using GSM8K training data, we evaluated two different options for the ground-truth answers to be used in training. The first was to use the original ground-truth answers provided with the dataset. The second was to instead use the baseline model’s greedy generations as the ground-truth answers, since the goal is to recover original model performance. To evaluate between the two options, we train with all scaling factors left free (no blocks removed) on both datasets. As we can see in Table 7, there seems to be a significant performance difference between the two options, with the model-generations-as-ground-truth scoring higher in both categories (flexible extract and strict match) for GSM8K. Additionally, we actually see improvement in strict match performance over the baseline when doing model-generations-as-ground-truth training. We will see later as well that this holds even when layers are removed from the model.
@@
Figure 3: Correlation between average model-ground-truth cross-entropy loss given a removed layer and benchmark performance on GSM8K (flexible extract and strict match).
@@ A.3 Base Model Selection
It is important to note that the selection of Llama-3.1-8B-Instruct as a base model was intentional, as it serves as an example of a model where all layers meaningfully transform the hidden states and contribute to the model’s outputs, better showing the generalizability of Hopscotch. For example, we show that with a model like our fine-tuned Granite-3.0-8B, the process of selecting attention blocks to remove is trivial, as there are a number of non-contributing blocks in the model. We illustrate why it is possible to skip attention layers by analyzing the largest singular values of the weight matrices across different layers. Singular values quantify how much a transformation can stretch or distort the input. Specifically, for a linear transformation $\bm{x}\to\mathbf{W}\bm{x}$, the inequality $\|\mathbf{W}\bm{x}\|_{2}\leq\sigma_{\max}(\mathbf{W})\|\bm{x}\|_{2}$ implies that if $\sigma_{\max}(\mathbf{W})$ is small and $\bm{x}$ is bounded, the output of the transformation will be near zero. This suggests that skipping such transformations would have minimal effect on the model’s output.
We present results using two open-source LLMs: Granite-3.0-8B (Granite Team, 2024) and Llama-3.1-8B-Instruct (Grattafiori et al., 2024) in Figure 4. The Granite model contains layers with negligible singular values, suggesting limited contribution to the output. Llama-3.1-8B-Instruct generally has larger and more consistent singular values. When we remove layers from Llama, we see a non-negligible drop in benchmark performance. With Granite, however, we see in Table 8 that removing attention blocks from layers with maximum singular values of zero results in no loss in model performance for the case of GSM8K.
@@ A.4 Using Loss as an Approximation of Benchmark Performance
As seen in Figure 3, training loss when removing layers serves as strong indication of how the removal will affect overall model performance. Thus, at any given model state, we can know the best attention block to remove by finding the block that, when removed, provides the lowest loss. Since we simply need an indication via quick conversion, rather than an optimized loss, we can simply run our training method for a single epoch with an increased learning rate (1e-2), and take the average training loss as a comparable value for each block.

### v2

@@ 1 Introduction
Large language models (LLMs) continue to grow in size, driven by “scaling laws” that suggest larger models tend to yield better performance (Hestness et al., 2017; Hoffmann et al., 2022; Henighan et al., 2020). Adding more attention blocks increases model capacity, but self-attention is also the most expensive operation in LLMs. Unlike MLP blocks, attention incurs a quadratic computational cost with respect to sequence length, making it a dominant factor in inference-time efficiency. However, not all attention blocks are equally important for every task, and some may carry redundant information. In this paper, we explore whether entire attention blocks can be skipped without significant performance degradation.
@@
We evaluate Hopscotch on instruction-tuned models, including Llama-3.1-8B-Instruct and Qwen2.5-7B-Instruct, and find that it can successfully skip up to 7 attention blocks with less than a 3% average performance drop across diverse tasks while yielding up to 15% inference speedup (see Section 4.6). With 4 blocks removed, we retain over 98.6% of baseline accuracy. These results highlight structural redundancy in attention blocks. We measure distributional shift in hidden representations and find that Hopscotch significantly reduces deviation from the original model, compared to unscaled attention block skipping. Finally, we show that Hopscotch is compatible with leading quantization techniques, such as GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2024).
@@ 2 Related Work
Skipping computation in LLMs. Several recent studies have explored the possibility of reducing computation in LLMs by selectively skipping parts of the model. Shukor and Cord (2024) show that in multimodal LLMs, certain attention blocks contribute less to performance in vision-language tasks and can be skipped in these cases with minimal degradation. Further investigations on the true impact of attention blocks at different depths (Ben-Artzy and Schwartz, 2024; He et al., 2024) find that the effect is highly block-specific. Our work builds on these insights by providing a principled mechanism to identify and compensate for less informative attention blocks using learned scaling.
@@ 3 The Hopscotch method
Method gsm8k FE gsm8k SM arc_challenge gpqa social_iqa pawsx
Baseline 83.10 66.94 64.33 29.07 49.85 63.26
4 Blocks Hopscotch 81.05 77.03 62.42 29.53 49.49 61.90
ShortGPT 62.33 58.45 57.00 28.44 47.59 59.49
FinerCut 75.80 61.00 53.58 27.77 46.93 54.98
7 Blocks Hopscotch 79.38 75.82 61.17 29.39 48.93 60.44
ShortGPT 1.90 1.29 46.42 27.35 43.76 57.62
FinerCut 42.50 6.80 52.65 28.10 46.47 54.87
10 Blocks Hopscotch 67.93 62.58 57.80 29.83 48.23 61.06
Table 1: Accuracy (%) of Llama-3.1-8B-Instruct on various benchmarks after skipping attention blocks (or full decoder blocks for ShortGPT). The table shows the scores of the original model compared to Hopscotch as well as prior methods when skipping 4, 7, or 10 blocks.
@@ 4.1 Benchmarks and Setup
Figure 2: A plot of the single-epoch average loss per block removed during step one of Hopscotch for Llama.
Figure 1 compares the benchmark performance of our full method, Hopscotch layer selection with scaling, against four baselines: (i) the original unmodified model (baseline), (ii) random block removal with no scaling, (iii) random block removal with Hopscotch scaling, and (iv) Hopscotch-selected layers with no scaling. When simply removing sets of arbitrary attention blocks near the center of the model, performance remained within a few points for all but GSM8K, which degraded drastically. When using Hopscotch scaling, even without layer selection, we were able to recover a significant portion of the GSM8K performance, while also seeing slight improvement in other domains. For this reason, we use the math domain and GSM8K as our primary test case and training recovery set for the Hopscotch method, while continuing to evaluate on the full set of benchmarks.
@@
We also compare directly to two leading prior works: SparseGPT Frantar and Alistarh (2023) and FinerCut Zhang et al. (2024). Hopscotch consistently outperforms both, especially on challenging tasks such as GSM8K, where other methods degrade significantly after removing four blocks and fail entirely beyond seven. Against ShortGPT, a key note is that the method removes entire layers, rather than keeping MLP blocks intact. This means for a fair comparison in terms of inference-time efficiency, one should compare ShortGPT with four layers removed to Hopscotch with at least six attention blocks removed (see Section 4.6 on how attention blocks contribute to inference time). We show that even with seven blocks removed, Hopscotch still outperforms across the board. For a closer comparison in terms of parameters removed, with 10 attention blocks removed, Hopscotch corresponds to about 83.3% of the parameter reduction by ShortGPT with 4 layers, yet achieves a relative inference speedup of $\mathbf{165.06\%}$. In this setting, Hopscotch still outperforms ShortGPT with 4 layers across all benchmarks (see Table 1). FinerCut also only removes attention blocks, making comparisons balanced with equal block removal. The primary remaining differences lie in the method of candidate block selection, as well as the critical discovery of the importance of scaling parameters.
@@
Table 2: Qwen2.5-7B-Instruct performance (accuracies) after skipping 4 attention blocks via Hopscotch.
@@ 4.5 Compatibility with Quantization
Baseline (Instruct) 67.0 83.9
GPTQ 68.0 83.9
GPTQ + Hopscotch (4) 77.3 79.4
GPTQ + Hopscotch (7) 75.1 78.5
Hopscotch (4) + GPTQ 77.9 79.2
Hopscotch (7) + GPTQ 73.8 78.5
AWQ 65.0 81.2
AWQ + Hopscotch (4) 71.4 75.1
AWQ + Hopscotch (7) 69.5 73.8
Hopscotch (4) + AWQ 75.2 78.3
Hopscotch (7) + AWQ 73.1 79.2
@@ 4.6 Effects on Model Efficiency
In a given forward pass for Llama-3.1-8B-Instruct, we experimentally find that $\sim$66% of time is spent in decoder attention blocks. In a model with 32 hidden layers, this means $\sim$2.06% of the forward pass is spent in each block. For example, for a query with an average forward pass time of 0.054 seconds, the time spent in a single attention block is on average 0.0011 seconds. The effect of removing a subset of attention blocks on overall inference time is summarized in Table 5.
Attn. Blocks
Removed
Inference
Time Reduction
@@
Given that Llama-3.1-8B-Instruct has 32 decoder layers and a total of $\sim$8B parameters, each layer contains approximately 250250M parameters. Removing one attention block saves roughly 8383M parameters, which is about $1.04\%$ of the model. Removing seven attention blocks yields a total reduction of $7\times 83$M = 581581M parameters, or about $7.28\%$ of the total model size. Since each parameter occupies 2 bytes in float16, we can estimate the memory savings accordingly (Table 6).
Attn. Blocks
Removed
Parameters
Reduction
Memory
Reduction
@@
[…]
@@ A.1 Data and Loss Function Selection
Using GSM8K training data, we evaluated two different options for the ground-truth answers to be used in training. The first was to use the original ground-truth answers provided with the dataset. The second was to instead use the baseline model’s greedy generations as the ground-truth answers, since the goal is to recover original model performance. To evaluate between the two options, we train with all scaling factors left free (no blocks removed) on both datasets. As we can see in Table 7, there seems to be a significant performance difference between the two options, with the model-generations-as-ground-truth scoring higher in both categories (flexible extract and strict match) for GSM8K. Additionally, we actually see improvement in strict match performance over the baseline when doing model-generations-as-ground-truth training. We will see later as well that this holds even when blocks are removed from the model.
@@
Figure 3: Correlation between average model-ground-truth cross-entropy loss given a layer with attention block removed and benchmark performance on GSM8K (flexible extract and strict match).
@@ A.3 Base Model Selection
It is important to note that the selection of Llama-3.1-8B-Instruct as a base model was intentional, as it serves as an example of a model where all layers meaningfully transform the hidden states and contribute to the model’s outputs, better showing the generalizability of Hopscotch. For example, we show that with a model like our fine-tuned Granite-3.0-8B, the process of selecting attention blocks to remove is trivial, as there are a number of non-contributing blocks in the model. We illustrate why it is possible to skip attention blocks by analyzing the largest singular values of the weight matrices across different layers. Singular values quantify how much a transformation can stretch or distort the input. Specifically, for a linear transformation $\bm{x}\to\mathbf{W}\bm{x}$, the inequality $\|\mathbf{W}\bm{x}\|_{2}\leq\sigma_{\max}(\mathbf{W})\|\bm{x}\|_{2}$ implies that if $\sigma_{\max}(\mathbf{W})$ is small and $\bm{x}$ is bounded, the output of the transformation will be near zero. This suggests that skipping such transformations would have minimal effect on the model’s output.
We present results using two open-source LLMs: Granite-3.0-8B (Granite Team, 2024) and Llama-3.1-8B-Instruct (Grattafiori et al., 2024) in Figure 4. The Granite model contains layers with negligible singular values, suggesting limited contribution to the output. Llama-3.1-8B-Instruct generally has larger and more consistent singular values. When we remove blocks from Llama, we see a non-negligible drop in benchmark performance. With Granite, however, we see in Table 8 that removing attention blocks from layers with maximum singular values of zero results in no loss in model performance for the case of GSM8K.
@@ A.4 Using Loss as an Approximation of Benchmark Performance
As seen in Figure 3, training loss when removing blocks serves as strong indication of how the removal will affect overall model performance. Thus, at any given model state, we can know the best attention block to remove by finding the block that, when removed, provides the lowest loss. Since we simply need an indication via quick conversion, rather than an optimized loss, we can simply run our training method for a single epoch with an increased learning rate (1e-2), and take the average training loss as a comparable value for each block.